### v4

@@ E.3   Decomposing the SLATET
In the previous subsection, we investigate how dCDH’s identifying assumptions ex ante restrict the treatment adoption behavior across units between exposed and unexposed groups. In this subsection, under these restrictions, we show that dCDH’s target parameter, the SLATET, can be decomposed into a weighted average of two different causal parameters.
The following theorem presents the decomposition result for the SLATET under Fuzzy DID designs, and clarifies the difference between dCDH’s target parameter (SLATET) and ours (LATET).
@@
Suppose Assumptions 24 - 28 hold. Then, we can decompose the SLATET into two different causal parameters:
@@
Theorem 7 shows that we can decompose the SLATET into a weighted average of two different causal parameters under Fuzzy DID designs: one parameter measures the treatment effects among the time compliers $CM^{T}$ in an exposed group and the other measures the treatment effects among the type $(CM^{Z}\land NT^{T})$ in an exposed group.
This theorem has two concerning implications. First, the interpretation of the SLATET is less clear than that of the LATET because it consists of two different causal parameters. Although the SLATET captures the ”local” average treatment effect, it may be difficult to interpret economically in a given application because one of the two causal parameters captures the treatment effects, for those who are induced to treatment by time.
Second, the SLATET may not be policy-relevant even when the researchers exploit the policy change of their interest as an instrument for treatment. Suppose that the instrument represents an important policy lever for researchers, who are intrinsically interested in the treatment effects among the compliers $CM^{Z}$ in an exposed group. In this situation, our target parameter, the LATET, is a policy-relevant parameter by construction. By contrast, the SLATET only captures the treatment effects among the type $(CM^{Z}\land NT^{T})$, which is the sub-population of the compliers $CM^{Z}$, even when the proportion of the time compliers $CM^{T}$ is zero within that group. As the share of the time compliers $CM^{T}$ increases, the SLATET puts more weight on the treatment effects among the time compliers, which contaminates the interpretation of the magnitude of this parameter even in such policy-relevant contexts.
Using the decomposition result for the SLATET, we finally clarify the interpretation of Assumption 29. When Assumption 25 (Stable percentage of treated units in an unexposed group) is violated, dCDH additionally consider Assumption 29 for the Wald-DID estimand to identify the SLATET, and call this the ”homogeneous” treatment effect assumption between exposed and unexposed groups.
However, we show that this assumption cannot be interpreted as requiring the ”homogeneous” treatment effect among the same sub-population between exposed and unexposed groups. If we do not assume Assumption 25, we have Table 11 in an unexposed group instead of Table 10 in Appendix E.2. Here, we assume that Assumption 26 and 31 exclude the defiers $DF^{Z}$ and the time defiers $DF^{T}$ in an unexposed group, respectively.
@@
By combining equations (42) and (43), we can conclude that Assumption 29 ($STATET=SLATET^{\prime}$) does not require the homogeneous treatment effects in the same sub-population between the two groups. Rather, this assumption requires that a weighted average of the treatment effects among the time compliers $CM^{T}$ and the type $(CM^{Z}\land NT^{T})$ in an exposed group should be equal to the treatment effect among the time compliers $CM^{T}$ in an unexposed group. This interpretation implies that the requirement of the ”homogeneous” treatment effect in Assumption 29 is less clear than what one might expect.
E.4   Revisiting the ”issue” regarding the use of the Wald-DID estimand
Finally, we revisit the ”issue” raised by dCDH regarding the use of the Wald-DID estimand in Fuzzy DID designs, and show that their argument is misguided. We also clarify why dCDH exclude so many types in exposed and unexposed groups as we see in Appendix E.2. Hereafter, we assume Assumptions 24 - 28 in Fuzzy DID designs.
dCDH argue that the Wald-DID estimand requires Assumption 28 (stable treatment effect over time) to identify the SLATET, and propose alternative estimands that do not rely on this assumption. We show that their argument stems from their reliance on the parallel trends assumption in the untreated outcome, which is common in DID designs, instead of on the parallel trends assumption in the unexposed outcome as in this article. Next, we argue that the former type of parallel trends assumption is unsuitable for DID-IV settings.
First, we clarify the interpretation of Assumption 28 by introducing the additional notation. This assumption is actually conditional on $D_{0}(Z)=1$ in addition to $Z$ and $T$, but dCDH do not provide the reason why this strong restriction is required only for the sub-population. For each time group $g^{t}$, let $w_{g^{t},z}=Pr(G^{T}=g^{t}|Z=z)$ denote the population share conditional on group $Z=z$, and let $w_{g^{t},z,D_{0}(0)=1}=Pr(G^{T}=g^{t}|Z=z,D_{0}(0)=1)$ denote the population share conditional on $D_{0}(0)=1$ in addition to group $Z=z$. Additionally, let $\delta_{z,t,g^{t}}=E[Y(1)-Y(0)|Z=z,T=t,G^{T}=g^{t}]$ denote the average treatment effect conditional on group $Z=z$, time $T=t$, and time group $G^{T}=g^{t}$.
Then, from Table 9 and Table 10, we can decompose Assumption 28 as follows:
$E[Y(1)-Y(0)|Z=0,T=1,D_{0}(0)=1]$ $=E[Y(1)-Y(0)|Z=0,T=0,D_{0}(0)=1]$
@@ E.4   Revisiting the ”issue” regarding the use of the Wald-DID estimand
Conditions (44) and (45) are required for Fuzzy DID designs because dCDH rely on the parallel trends assumption in the untreated outcomes. In DID-IV settings, we can not observe the average time trends of untreated potential outcomes for the time always takers $AT^{T}$ in an unexposed group and the time always takers $AT^{T}$ and the time defiers $DF^{T}$ in an exposed group because they always adopt the treatment during the two periods (see Table 9 and Table 10). Therefore, if we assume the parallel trends assumption in the untreated outcomes, we should recover them by using conditions (44) and (45).
To see this formally, we use our notation to briefly sketch the proof of dCDH on the Wald-DID estimand identifying SLATET under fuzzy DID designs. First, consider the conditional expectation $E[Y|Z=0,T=1]-E[Y|Z=0,T=0]$, which appears in the numerator of the Wald-DID estimand. We can decompose this quantity into two terms.
@@
Then, by noticing that the denominator of the Wald-DID estimand identifies $Pr(SW|Z=1)$ under Assumption 25 and Assumption 26, the Wald-DID estimand captures the SLATET under Assumption 27 (parallel trends assumption in the untreated outcome) and conditions (44) and (45). From the discussions, one can conclude that the parallel trends assumption in the untreated outcome leads dCDH to impose conditions (44) and (45).
@@
dCDH exclude the type $NT^{Z}\land DF^{T}$ in an exposed group and the time defiers $DF^{T}$ in an unexposed group because if we add and subtract the average time trends of the untreated outcome for these types, we additionally identify the treatment effects among them in period 00 (not in period $1$), as they adopt the treatment in period 00, but leave it in period $1$ (see Tables 9-10). This identification result does not seem intuitive because the identification variation comes from the policy shock (instrument) in period $1$.
@@
dCDH exclude the time compliers $CM^{T}$ in an unexposed group because if we add and subtract the average time trends of the untreated outcome for that type, we additionally capture the treatment effects among the time compliers $CM^{T}$ from the conditional expectation $E[Y|Z=0,T=1]-E[Y|Z=0,T=0]$, which implies that the SLATET in an exposed group is offset by the SLATET’ in an unexposed group.
We have confirmed that dCDH’s argument regarding the use of the Wald-DID estimand rests on the parallel trends assumption in the untreated outcome in DID designs. Finally, we argue that this assumption is unsuitable for DID-IV settings for three reasons.
First, this assumption is not sufficient to capture the average time trends of the outcome even in an unexposed group. This is because in DID-IV settings, the units are allowed to take the treatment in the absence of the instrument, and we cannot observe the untreated outcomes for these types in an unexposed group.
Second, in most of the DID-IV applications, we can not impute the untreated potential outcomes in general. For instance, Black et al. (2005) estimates the causal link between parents’ and children’s education attainment, exploiting variation in the timing of the implementation of the school reform across municipalities and cohorts as an instrument for parents’ education attainment. In the authors’ setting, the parallel trends assumption in the untreated outcome requires the data to contain parents with zero education attainment, which is unrealistic and not what the authors would have in mind in practice.
Finally, this assumption is generally not testable using the pre-exposed period data in a given application. This is apparent, as some units can already adopt the treatment before period 00, and we cannot compare the pre-trend of the untreated outcome between exposed and unexposed groups.
dCDH propose the placebo test in their appendix, but we show that it is not enough for checking the plausibility of the parallel trends assumption in the untreated outcome. dCDH’s test has two steps. First, we test whether the share of treated units is zero in both groups from $T=-1$ to $T=0$.
@@
$E[Y|Z=1,T=0]-E[Y|Z=1,T=-1]=E[Y|Z=0,T=0]-E[Y|Z=0,T=-1].$ (48)
Although dCDH ”assume” the conditions (46) and (47), they note that when one of the two conditions fails, the second test can ”no longer be used to test” Assumption 27. We, therefore, describe the dCDH’s test in two steps.
dCDH’s test has some drawbacks. First, the conditions (46) and (47), which should be checked in the first step, are strong, but dCDH do not explain the way ahead when these conditions are violated in practice. Second, in this test, the second step does not directly test Assumption 27 in the pre-exposed periods, as it does not compare the evolution of the mean untreated outcome between exposed and unexposed groups before period $T=0$. Indeed, dCDH describe the equality (48) as a necessary condition for Assumption 27 and Assumption 28 to be satisfied under the conditions (46) and (47). Finally, it is unclear how this test should be conducted with multiple pre-exposed period data. dCDH assume ”for instance that data is available for the period $T=-1$”, but they do not indicate how this test can be performed in multiple pre-exposed period settings.
@@
$E[D|Z=1,T=0]-E[D|Z=1,T=-1]=E[D|Z=0,T=0]-E[D|Z=0,T=-1].$ (49)
Under repeated cross section settings, the conditions (48) - (49) coincide with the conditions (3) -(4).

### v5

@@ E.3   Decomposing the SLATET
In this subsection, we show that dCDH’s target parameter, the SLATET, can be decomposed into a weighted average of two different causal parameters under Fuzzy DID designs.
The following theorem presents the decomposition result for the SLATET, and clarifies the difference between dCDH’s target parameter (SLATET) and ours (LATET).
@@
Suppose Assumptions 26-30 hold. Then, we can decompose the SLATET into two different causal parameters:
@@
Theorem 6 shows that the SLATET can be decomposed into a weighted average of two different causal parameters under Fuzzy DID designs: one parameter measures the treatment effects among the time compliers $CM^{T}$ in an exposed group and the other measures the treatment effects among the type $(CM^{Z}\land NT^{T})$ in an exposed group.
This theorem has a concerning implications; the SLATET may not be policy-relevant even when the instrument is based on the policy change of interest to the researcher. Suppose that the instrument represents an important policy lever for researchers, and they are intrinsically interested in the treatment effects among the compliers $CM^{Z}$ in an exposed group. In this setting, our target parameter, the LATET, is a policy-relevant by construction. However, the SLATET may not be policy-relevant because this parameter captures the treatment effects among the sub-population of the compliers and the time compliers, who are affected by time but not affected by the instrument. As the share of the time compliers $CM^{T}$ increases, the SLATET puts more weight on the treatment effects among the time compliers, which contaminates the interpretation of the magnitude of this paramter even in such policy-relevant contexts.
Using the decomposition result for the SLATET, we finally clarify the interpretation of Assumption 31. We show that this assumption cannot be interpreted as requiring the homogeneous treatment effect among the same sub-population between exposed and unexposed groups.
If we do not assume Assumption 27, we have Table 5 in an unexposed group. [^30] Here, we assume that Assumption 28 and 33 exclude the defiers $DF^{Z}$ and the time defiers $DF^{T}$ in an unexposed group, respectively.
@@
By combining equation (31) with (32), one can see that Assumption 31 ($STATET=SLATET^{\prime}$) does not require the homogeneous treatment effects in the same sub-population between the two groups. Rather, this assumption requires that a weighted average of the treatment effects among the time compliers $CM^{T}$ and the type $(CM^{Z}\land NT^{T})$ in an exposed group should be equal to the treatment effect among the time compliers $CM^{T}$ in an unexposed group. This interpretation implies that the requirement of the homogeneous treatment effect in Assumption 31 is less clear than what one might expect.
E.4   Revisiting the issue regarding the use of the Wald-DID estimand
Finally, we revisit the issue pointed by dCDH regarding the use of the Wald-DID estimand, and show that this problem is not serious than one might expect, and does not arise under our DID-IV designs. We also clarify why dCDH exclude so many types in exposed and unexposed groups as seen in Tables 3-4. Hereafter, we assume Assumptions 26-30 in Fuzzy DID designs.
dCDH argue that the Wald-DID estimand requires Assumption 30 to identify the SLATET, and propose alternative estimands that do not rely on this assumption. We first show that their argument depends on the parallel trends assumption in DID designs. Next, we argue that this assumption would be difficult to justify in DID-IV settings.
First, we clarify the interpretation of Assumption 30 by introducing the additional notation. This assumption is actually conditional on $D_{0}(Z)=1$ in addition to $Z$ and $T$, but it is not immediately clear from dCDH why this strong restriction is required only for the sub-population. For each time group $g^{t}$, let $w_{g^{t},z}=Pr(G^{T}=g^{t}|Z=z)$ denote the population share conditional on group $Z=z$, and let $w_{g^{t},z,D_{0}(0)=1}=Pr(G^{T}=g^{t}|Z=z,D_{0}(0)=1)$ denote the population share conditional on $D_{0}(0)=1$ in addition to group $Z=z$. Additionally, let $\delta_{z,t,g^{t}}=E[Y(1)-Y(0)|Z=z,T=t,G^{T}=g^{t}]$ denote the average treatment effect conditional on group $Z=z$, time $T=t$, and time group $G^{T}=g^{t}$.
Then, from Tables 3-4, we can decompose Assumption 30 as follows:
$E[Y(1)-Y(0)|Z=0,T=1,D_{0}(0)=1]$
$=$ $E[Y(1)-Y(0)|Z=0,T=0,D_{0}(0)=1]$
@@ E.4   Revisiting the issue regarding the use of the Wald-DID estimand
Conditions (33) and (34) are required for Fuzzy DID designs because dCDH rely on the parallel trends assumption in the untreated outcome. In DID-IV settings, we can not observe the average time trends of untreated potential outcomes for the time always takers $AT^{T}$ in an unexposed group and the time always takers $AT^{T}$ and the time defiers $DF^{T}$ in an exposed group because they always adopt the treatment during the two periods. Therefore, if we assume the parallel trends assumption in the untreated outcomes, we should recover them by using conditions (33) and (34).
To see this formally, we use our notation to briefly sketch dCDH’s proof about how the Wald-DID estimand identifies the SLATET under fuzzy DID designs. First, consider the conditional expectation $E[Y|Z=0,T=1]-E[Y|Z=0,T=0]$, which appears in the numerator of the Wald-DID estimand. We can decompose this quantity into two terms.
@@
Then, by noticing that the denominator of the Wald-DID estimand identifies $Pr(SW|Z=1)$ under Assumption 27 and Assumption 28, the Wald-DID estimand captures the SLATET under Assumption 29 (parallel trends assumption in the untreated outcome) and conditions (33) and (34). From the above discussion, one can see that imposing the parallel trends assumption in the untreated outcome leads dCDH to assume conditions (33) and (34).
@@
dCDH exclude the type $NT^{Z}\land DF^{T}$ in an exposed group and the time defiers $DF^{T}$ in an unexposed group because if we add and subtract the average time trends of the untreated outcome for these types, we additionally identify the treatment effects among them in period 00 (not in period $1$), as they adopt the treatment in period 00, but leave it in period $1$ (see Tables 3-4). This identification result, however, is not intuitive because the identification variation comes from a policy shock (instrument) in period $1$.
@@
dCDH exclude the time compliers $CM^{T}$ in an unexposed group in order to avoid the SLATET in an exposed group being offset by the SLATET’ in an unexposed group. If we add and subtract the average time trends of the untreated outcome for the time compliers, we additionally capture the SLATET’ in an unexposed group from the conditional expectation $E[Y|Z=0,T=1]-E[Y|Z=0,T=0]$.
Thus far, we have confirmed that dCDH’s argument rests on the parallel trends assumption in the untreated outcome. Finally, we argue that this assumption would be difficult to rely on in DID-IV settings for three reasons.
[…]
Finally, this assumption is generally not testable using the pre-exposed period data, as some units can already adopt the treatment before period 00, and we cannot compare the pre-trend of the untreated outcome between exposed and unexposed groups.
Although dCDH propose the placebo test in their appendix, it has shortcomings for checking the plausibility of the parallel trends assumption in the untreated outcome. dCDH’s test has two steps. [^31] Although dCDH ”assume” the conditions (35) and (36), they note that when one of the two conditions fails, the second test can ”no longer be used to test” Assumption 29. We, therefore, describe the dCDH’s test in two steps. First, we test whether the share of treated units is zero in both groups from $T=-1$ to $T=0$.
@@
$E[Y|Z=1,T=0]-E[Y|Z=1,T=-1]$ (37)
$=$ $E[Y|Z=0,T=0]-E[Y|Z=0,T=-1].$ (38)
dCDH’s test has some weaknesses. First, the conditions (35) and (36), which should be checked in the first step, are strong, but no guidance is provided when these conditions are violated in practice. Second, in this test, the second step does not directly test Assumption 29 in the pre-exposed periods, as it does not compare the evolution of the mean untreated outcome between exposed and unexposed groups before period $T=0$. Indeed, dCDH describe the equality (37) as a necessary condition for Assumption 29 and Assumption 30 to be satisfied under the conditions (35) and (36). Finally, it is not clear how this test should be conducted with multiple pre-exposed period data. Although dCDH assume “for instance that data is available for the period $T=-1$”, it is not indicated how this test can be performed in multiple pre-exposed period settings.
@@
$E[D|Z=1,T=0]-E[D|Z=1,T=-1]$ (39)
$=$ $E[D|Z=0,T=0]-E[D|Z=0,T=-1].$ (40)
Under repeated cross section settings, the conditions (37)-(39) coincide with the conditions (1)-(2). This indicates that the rationale behind dCDH’s test is more suitable for the assessment of parallel trends assumptions in DID-IV designs.
E.5   Other connections
In this subsection, we document the relationship between DID-IV and Fuzzy DID under non-binary, ordered treatment, and multiple groups settings. We also compare the Wald-DID estimand to the time-corrected Wald estimand considered in dCDH.
E.5.1 Non-binary, ordered treatment
In this subsection, we investigate the relationship between DID-IV to Fuzzy DID under non-binary, ordered treatment settings.
First, we compare the target parameter between the two designs. With a non-binary, ordered treatment $D\in\{0,1,\dots,J\}$, dCDH consider the following target parameter, which we call the switcher average causal response (SACR).
$SACR\equiv\sum_{j=1}^{J}w_{j}\cdot E(Y(j)-Y(j-1)|Z=1,T=1,D_{0}(0)<j<D_{1}(1)),$
where $w_{j}$ is
$w_{j}=\frac{P(D\geq j|Z=1,T=1)-P(D\geq j|Z=1,T=0)}{E[D|Z=1,T=1]-E[D|Z=1,T=0]}.$
[…]
Next, we compare the identifying assumptions between the two designs. In order for the Wald-DID estimand to identify the SACR, dCDH replace Assumption 28 with the following one.
Assumption 34 (Ordered treatment equation).
$D=\sum_{j=1}^{J}\mathbf{1}\{V\geq v^{j}_{ZT}\},\hskip 5.69054pt\text{with}\hskip 5.69054pt-\infty=v^{0}_{ZT}<v^{1}_{ZT}\dots<v^{J+1}_{ZT}=\infty\hskip 5.69054pt\text{and}\hskip 5.69054ptV\mathop{\perp\!\!\!\!\perp}T|Z.$
In addition, dCDH replace the condition (27) in Assumption 26 with the condition (41), and Assumption 27 with the condition (42), respectively.
$D_{11}\succeq D_{10},$ (41)
$D_{01}\sim D_{00},$ (42)
where $D_{ZT}$ denotes the treatment conditional on $Z$ and $T$, $\succeq$ denotes stochastic dominance between two random variables, and $\sim$ denotes equality in distribution.
Note that the condition (42) is stronger than Assumption 27 because this assumption requires that the distribution of the treatment should be stable during the two periods in an unexposed group.
In our DID-IV designs, the identifying assumptions required for the Wald-DID estimand to capture the ACRT are unchanged under non-binary, ordered treatment settings. In addition, Assumptions 17-18 are indirectly testable using the pre-exposed period data in practice.
E.5.2 Multiple groups
In multiple groups settings, our DID-IV design is not directly compatible to Fuzzy DID design because dCDH exploit the group structure in the data ($G\in\{0,1,\dots,\bar{g}\}$) for the Wald-DID estimand to identify the following target parameter.
$\Delta^{\ast}=E[Y(1)-Y(0)|\cup_{g=0}^{\bar{g}}S_{g},T=1],$
where $S_{g}=\{D_{0}(g)\neq D_{1}(g),G=g\}$ represents the switchers in group $G=g$ [^32] Following $D_{t}(Z)$, we define $D_{t}(G)=\mathbf{1}\{V\geq v_{Gt}\}$ to be the treatment status under time $T=t$ in group $G$.
[…]
In some applications, a researcher might prefer our approach to dCDH’s approach for three reasons. First, our approach yields a policy-relevant parameter by construction if the instrument represents a policy change of interest to the researcher. Second, in multiple period settings, our approach avoids the need to estimate the super groups in each consecutive period $(t-1,t)$, which simplifies both estimation and inference, making our method easier to implement in practice.
Finally, in our approach, the monotonicity assumption can be easy to justify if the instrument is based on a policy shock. In dCDH’s approach, this assumption might be harder to defend in practice because the group variable $G$ typically represents the geographic feature in the data such as municipality, state, or country. [^34] Note that, in multiple groups settings, dCDH also maintain the monotonicity assumption on each group pair $(g,g^{\prime})$ through Assumption 28.
Remark E.4.3 (Multiple groups settings with an ordered treatment).
[…]
E.5.3 Alternative estimand
In this subsection, we compare the Wald-DID estimand to the alternative estimand proposed by dCDH that does not rely on Assumption 29 and 30.353535In this subsection, we do not consider Wald-CIC estimand proposed by dCDH, as we focus on the mean treatment effect in this article.
As an alternative to the Wald-DID estimand, dCDH propose the time-corrected Wald (TC-Wald) estimand defined below.
$W_{TC}\equiv\frac{E[Y|Z=1,T=1]-E[Y+(1-D)\delta_{0}+D\delta_{1}|Z=1,T=0]}{E[D|Z=1,T=1]-E[D|Z=1,T=0]},$
where $\delta_{d}=E[Y|Z=0,T=1,D_{1}(0)=d]-E[Y|Z=0,T=0,D_{0}(0)=d]$.
In order for the TC-Wald estimand to identify the SLATET, dCDH consider the following identifying assumption instead of imposing Assumption 29 and 30.
Assumption 35 (Conditional common trends).
For all $d\in\mathcal{S}(D),$
$E[Y(d)|Z=1,T=1,D_{0}(0)=d]-E[Y(d)|Z=1,T=1,D_{0}(0)=d]$
$=$ $E[Y(d)|Z=0,T=1,D_{0}(0)=d]-E[Y(d)|Z=0,T=1,D_{0}(0)=d].$
Note that with a binary treatment, this assumption exploits the observed outcome path in an unexposed groups similar to Assumption 18 in our DID-IV designs (see Table 4).
While the target parameter differs between the two designs, we see that using the Wald-DID estimand under DID-IV designs would be more appealing than using the TC-Wald estimand under Fuzzy DID designs for two reasons. First, when using the Wald-DID estimand in DID-IV designs, researchers can easily assess the plausibility of parallel trends assumptions in practice, employing the pre-trends tests described in Section 5.3. Although the TC-Wald estimand avoids the stable treatment effect assumption under Fuzzy DID designs, it presents challenges in testing the plausibility of Assumption 35 in pre-exposed period data: it is conditional on $D_{0}(0)=d$ and imposes restrictions on the mean evolution of $Y(d)$ for all $d\in\mathcal{S}(D)$.
[…]